# Arabizi vs LLMs: Can the Genie Understand the Language of Aladdin?

### v1

@@ Abstract
In this era of rapid technological advancements, communication continues to evolve as new linguistic phenomena emerge. Among these is Arabizi, a hybrid form of Arabic that incorporates Latin characters and numbers to represent the spoken dialects of Arab communities. Arabizi is widely used on social media and allows people to communicate in an informal and dynamic way, but it poses significant challenges for machine translation due to its lack of formal structure and deeply embedded cultural nuances. This case study arises from a growing need to translate Arabizi for gisting purpose. It evaluates the capacity of different LLMs to decode and translate Arabizi, focusing on multiple Arabic dialects that have rarely been studied up until now. Using a combination of human evaluators and automatic metrics, this research project investigates the model’s performance in translating Arabizi into both Modern Standard Arabic and English. Key questions explored include which dialects are translated most effectively and whether translations into English surpass those into Arabic.
@@
Perla Al Almaoui1 Pierrette Bouillon1 Simon Hengchen1,2 1Faculté de traduction et d’interprétation, Université de Genève 2iguanodon.ai Correspondence: almaoui.perla@outlook.com
@@ 1 Introduction
Although there are approximately 420 million Arabic speakers worldwide, an intriguing linguistic paradox emerges: Modern Standard Arabic (MSA), the standardized form of the language, is the mother tongue of none. Instead, Arabs communicate through their regional dialects, which are vibrant linguistic hybrids influenced by Arabic and the historical languages of each region. These dialects have been honed by geographic, cultural, and historical factors, and can vary significantly even within a single country, resulting in a mosaic of over 60 distinct varieties. Arabizi (a fusion of "Arabic" and Englizi, the Arabic word for English) is an informal, non-standard writing system that emerged in the 1990s when Arabic keyboards were not widely available. It uses Latin characters and numbers, combining both transliteration and transcription mappings. Primarily used in online communication—such as short messages and comments on social media—Arabizi varies significantly across dialects and even within the same dialect Harrat et al. (2019).
@@
Despite its widespread use across digital platforms, and the recent focus on informal language and low-resource languages, Arabizi remains an unexplored area in natural language processing (NLP). It poses particular challenges due to its colloquial nature, variation across dialects and lack of standardization, as well as the scarcity of digital resources. In NLP, research on Arabizi has mainly focused on transliteration into Arabic (deromanization) at the character or word level, using different approaches Guellil et al. (2017); Shazal et al. (2020). Some studies explore interlinguistic machine translation (MT) techniques to and from Arabizi, employing various architectures and pipelines, mainly between English and Egyptian dialects (see Harrat et al. (2019) for a summary up until 2017). While some recent datasets for low resource language translation include romanization, they are not specifically focused on Arabizi. Flores benchmark Goyal et al. (2022), for example, is limited to the romanized transcription of MSA or Arabic dialects in Arabic scripts. The TerjamaBench dataset Momayiz et al. (2024) is an exception and includes entries in Darija, the Moroccan Arabic dialect, written in both Latin alphabet (Arabizi) and Arabic script, and their corresponding English translations.
Since there is a growing need to translate Arabizi into resource-rich languages on social media and other digital platforms, we conducted a case study to evaluate the feasibility of using large language models (LLMs) for out-of-the-box machine translation. The project began when the Brussels-based language technology company iguano-don.ai received a request from a client who wanted to know if short Arabizi texts could be translated for gisting purposes. The study involves a collaboration between the start-up and a professional translator with previous experience in Arabizi. Our contribution includes an authentic dataset in Arabizi for three dialects and a comparative evaluation of five LLMs using different prompting strategies. To our knowledge, this is the first study that has explored the direct translation from different dialects in Arabizi to MSA or English without prior deromanization.
@@ 2.1 Dialects
We decided to focus on the translation of three Arabic dialects from three distinct countries—Lebanon, Egypt, and Algeria—into two target languages: MSA, a less-resourced language, and English. These three countries were selected because their dialects represent distinct linguistic varieties. The Lebanese dialect aligns with the Levantine group and the Algerian dialect with the Maghrebi group, while the Egyptian dialect is exceptionally prominent due to its widespread use and cultural influence.
@@
Egyptian Arabic evolved through layers of historical migrations, demographic shifts, and ancient linguistic roots. It was heavily influenced by Coptic, the language of ancient Egypt, and later by Arabic after the Islamic conquest in the 7th century. Over time, Egyptian Arabic absorbed linguistic elements from Greek, Turkish, Italian, French, and English during various periods of occupation and cultural exchange. Regional variations within Egypt further enrich its linguistic diversity: northern regions, including the Delta and Cairo, feature subtle dialectal differences, while Upper Egypt’s Sa’idi Arabic retains more conservative features. Additionally, Bedouin communities in the Western Desert speak Arabic varieties that are distinct from urban Egyptian Arabic Souag (2009); Bettega et al. (2022).
@@ 2.2 Participants
After the corpus was collected, it was anonymized to ensure privacy. Subsequently, a professional Arabic-speaking translator translated the corpus into MSA and English, with these translations serving as reference texts for automatic metrics. The translation into Arabic represents an intralingual transformation from a dialectal and informal variety of Arabic to a formal and standardized form.
Table 1 presents the collected corpora, including the number of segments and tokens, the average number of tokens per sentence and the percentage of foreign and mixed words (code-switching).
@@ 3 Experimental Setup
  • model $\in$ [GPT4o, Llama 3, Claude, Gemini, Gemma, Mistral, Jais].
All models were prompted with a temperature of $0.5$. A discussion of the chosen models and prompts is available in 3.1 and in 3.2. Evaluation metrics are presented in 3.3. Our code will be shared upon acceptance.
@@ 3.1 Models
The models used in the experiments are all decoder-only transformer Vaswani et al. (2017) models generally called “generative LLMs”. We used a range of instruction-tuned LLMs of different parameter sizes (from 27B for Gemma to at least 70B for Llama3, while proprietary models are expected to be much larger) to cover various models, from open weights to proprietary, general purpose or, in the case of Jais, ones that specifically target the English-Arabic pair Sengupta et al. (2023).
In this paper, “Llama” refers to Llama 3.3 70B-Instruct (Dubey et al., 2024), “GPT-4o” to gpt-4o-2024-08-06 (OpenAI, 2024), [^2] https://platform.openai.com/docs/models/gpt-4o#gpt-4o “Claude” to Anthropic ’s 3.5 Sonnet,333anthropic.claude-3-5-sonnet-20240620-v1:0 “Gemma” to gemma-2-27b-it Gemma Team et al. (2024), and, also from Google, “Gemini” to the latest Gemini 1.5 Pro version Gemini Team et al. (2024). [^4] December 2024 release. “Mistral” is Mistral Large 24.11 from the eponymous company and, finally, “Jais” refers to jais-family-30b-16k-chat Sengupta et al. (2023).
@@ 3.2 Prompts
In order to achieve the best translation results, we built on He (2024)’s findings by assigning the role of a professional translator to our LLM. This approach outperformed both simpler prompts and those with excessive context. Furthermore, for each of the three main dialects, we used three prompt strategies: no-shot, one-shot, and two-shot, all written in English. These prompts were the same across regions, except for the specific mention of each dialect in the corresponding prompts. The examples used in the one- and two-shot configurations are not part of the evaluated set, and are from the Algerian and Lebanese dialects. We further refined and duplicated these prompt variations to cover two target languages: one set asked for translation into English and the other into MSA. Finally, all prompts were translated into Arabic by a native Arabic speaker who is also a professional translator. In total, we ran experiments with 36 unique prompts (3 regions * 3 strategies * 2 target languages * 2 prompt languages), or 18 per target language, which we used on all models. The prompts in English are available in Appendix A, while their equivalents in Arabic will be released as part of the code release upon acceptance.
@@ 3.3.2 Human Evaluation
For human evaluation, we adopted the direct assessment method, which evaluates translations based on a Likert scale ranging from 1 to 5 (higher is better) according to two key criteria: fluency and adequacy (See Appendix B). Due to time and human resource constraints, we did not manually annotate all translations. We instead sampled a random machine translation for both target languages and for each source sentence of our dataset. These machine translation outputs were sampled across all our variables, i.e. models, prompt languages, and prompt strategies. The resulting set, consisting of 671 segments (268 for Lebanon, 159 for Egypt and 244 for Algeria, see Table 1) for each target language, was then manually rated by two native speakers of Arabic who are professional translators, one of them being the first author of this study and the original translator of the dataset. We then calculated Cohen (1960) $\kappa$ to measure their agreement in terms of fluency and accuracy (see Table 2 for results). Cohen (1960) $\kappa$ results indicate moderate agreement for adequacy and lower agreement for fluency, with some variations across language pairs.
The set, which not only consisted of the reference and machine translation but also of the source sentence, was then iteratively fed into gpt4o in an “LLM-as-a-judge” setting (Zheng et al., 2023), with a prompt in English tasking the LLM to follow the human annotation guidelines. [^6] The prompt is shared in Appendix C. The LLM showed strong correlation with both human annotators, with Spearman (1904)’s $\rho$s comprised in the range from $0.457$ (annotator 2, fluency, Egypt to EN) to $0.844$ (annotator 2, adequacy, Egypt to AR). These results indicate that an LLM can be used as an easy way to gauge translation quality during model development. The different correlations as well as all the data for direct assessment can be found in Appendix E.
LB-AR LB-EN EG-AR EG-EN ALG-AR ALG-EN
Adequacy 0.56 0.39 0.34 0.33 0.49 0.45
Fluency 0.24 0.31 0.21 0.52 0.28 0.31
Table 2: Agreement scores (Cohen (1960)’s $\kappa$) for different language pairs for Adequacy and Fluency Koehn and Monz (2006). LB = Lebanon, EG = Egypt, ALG = Algeria, {AR, EN} = Translation to {Arabic, English}
@@ 4.1 Qualitative Error Analysis
Due to space constraints, this section will only provide some examples of the main errors (full annotation will be provided upon acceptance). Most models tend to mistranslate, especially when figurative language is used. A larger issue lies with Llama3 which tends to output words in another script when translating to MSA. An obvious example is the translation of the segment Almatar da (\<هذا المطار>, “this airport”) as \<المطر> да – transforming the “da” in Arabic to Cyrillic. The problem is not limited to Cyrillic, as characters in Latin and Chinese scripts can also be found in the output. Another type of failure specific to a model is Jais’ re-occuring hallucinations. The model often associates feelings of anger to an otherwise neutral message, leading to translations that are irrelevant and contain violent information.
@@ Appendix D Metrics Signatures
The BLEU signature is nrefs:1|case:mixed|eff:no|tok:13a|smooth:exp|version:2.5.1.
The chrF signature is nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.5.1.
The TER signature is nrefs:1|case:lc|tok:tercom|norm:no|punct:yes|asian:no|version:2.5.1.
@@ Appendix F Automatic Evaluation Results for all Models
Due to space constraints, we are presenting the results averaged over the two prompt languages. The full set will be shared upon acceptance.
@@
gpt4o EN no-shot 17.386 43.081 77.038 0.478 0.434 0.733
gpt4o EN one-shot 20.158 45.232 70.287 0.529 0.439 0.757
gpt4o EN two-shot 19.771 45.496 70.294 0.521 0.439 0.755
gpt4o AR no-shot 8.395 36.637 86.501 0.558 0.422 0.757
gpt4o AR one-shot 10.099 38.221 78.701 0.586 0.423 0.776
gpt4o AR two-shot 10.150 38.090 79.775 0.585 0.421 0.774

### v2

@@ Abstract
In this era of rapid technological advancements, communication continues to evolve as new linguistic phenomena emerge. Among these is Arabizi, a hybrid form of Arabic that incorporates Latin characters and numbers to represent the spoken dialects of Arab communities. Arabizi is widely used on social media and allows people to communicate in an informal and dynamic way, but it poses significant challenges for machine translation due to its lack of formal structure and deeply embedded cultural nuances. This case study arises from a growing need to translate Arabizi for gisting purposes. It evaluates the capacity of different LLMs to decode and translate Arabizi, focusing on multiple Arabic dialects that have rarely been studied up until now. Using a combination of human evaluators and automatic metrics, this research project investigates the models’ performance in translating Arabizi into both Modern Standard Arabic and English. Key questions explored include which dialects are translated most effectively and whether translations into English surpass those into Arabic.
@@
Perla Al Almaoui Faculté de traduction et d’interprétation Université de Genève almaoui.perla@outlook.com                        Pierrette Bouillon Faculté de traduction et d’interprétation Université de Genève pierrette.bouillon@unige.ch
Simon Hengchen Faculté de traduction et d’interprétation & iguanodon.ai Université de Genève simon.hengchen@unige.ch
footnotetext: ©  2025 The authors. This article is licensed under a Creative Commons 4.0 licence, no derivative works, attribution, CC-BY-ND.
@@ 1 Introduction
Although there are approximately 420 million Arabic speakers worldwide, an intriguing linguistic paradox emerges: Modern Standard Arabic (MSA), the standardized form of the language, is the mother tongue of none. Instead, Arabs communicate through their regional dialects, which are vibrant linguistic hybrids influenced by Arabic and the historical languages of each region. These dialects have been honed by geographic, cultural, and historical factors, and can vary significantly even within a single country, resulting in a mosaic of over 60 distinct varieties. Arabizi (a fusion of "Arabic" and Englizi, the Arabic word for English) is an informal, non-standard writing system that emerged in the 1990s when Arabic keyboards were not widely available. It uses Latin characters and numbers, combining both transliteration and transcription mappings. Primarily used in online communication—such as short messages and comments on social media—Arabizi varies significantly across dialects and even within the same dialect Harrat et al. (2019). For instance, the transcription of the following sentence in MSA \<أريد أن أكلّمك بموضوع> (I want to talk to you about something) in Arabizi can be "badde e7kik bi mawdu3" in Levantine Arabic, or "rani hab nahdar m3ak f wahd sujet" in Algerian Arabic.
@@
Despite its widespread use across digital platforms, and the recent focus on informal language and low-resource languages, Arabizi remains an unexplored area in natural language processing (NLP). It poses particular challenges due to its colloquial nature, variation across dialects and lack of standardization, as well as the scarcity of digital resources. In NLP, research on Arabizi has mainly focused on transliteration into Arabic (deromanization) at the character or word level, using different approaches Guellil et al. (2017); Shazal et al. (2020), and on the creation of a parallel annotated corpus of SMS and chat messages written in Arabizi and their corresponding Arabic script transliterations by Bies et al. (2014). Some studies explore interlinguistic machine translation (MT) techniques to and from Arabizi, employing various architectures and pipelines, mainly between English and Egyptian dialects (see Harrat et al. (2019) for a summary up until 2017). While some recent datasets for low resource language translation include romanization, they are not specifically focused on Arabizi. Flores benchmark Goyal et al. (2022), for example, is limited to the romanized transcription of MSA or Arabic dialects in Arabic scripts. The TerjamaBench dataset Momayiz et al. (2024) is an exception and includes entries in Darija, the Moroccan Arabic dialect, written in both Latin alphabet (Arabizi) and Arabic script, and their corresponding English translations.
Since there is a growing need to translate Arabizi into resource-rich languages on social media and other digital platforms, we conducted a case study to evaluate the feasibility of using large language models (LLMs) for out-of-the-box machine translation. The project began when the language technology company iguanodon.ai received a request from a client who wanted to know if short Arabizi texts could be translated for gisting purposes. The study involves a collaboration between the start-up and a professional translator with previous experience in Arabizi. Our contribution includes AladdinBench [Uncaptioned image], an authentic dataset in Arabizi for three dialects publicly available on huggingface [^1] https://huggingface.co/datasets/palmaoui/AladdinBench and a comparative evaluation of five LLMs using different prompting strategies.
To our knowledge, this is the first study that has explored the direct translation from different dialects in Arabizi to MSA or English without prior deromanization.
@@ 2.1 Dialects
We decided to focus on the translation of three Arabic dialects from three distinct countries—Lebanon, Egypt, and Algeria—into two target languages: MSA, a less-resourced language, and English. These three countries were selected because their dialects represent distinct linguistic varieties. The Lebanese dialect aligns with the Levantine group and the Algerian dialect with the Maghrebi group, while the Egyptian dialect is exceptionally prominent due to its widespread use and cultural influence Ayoubi (2022).
@@
Egyptian Arabic evolved through layers of historical migrations, demographic shifts, and ancient linguistic roots. It was heavily influenced by Coptic, the language of ancient Egypt, and later by Arabic after the Islamic conquest in the 7th century. Over time, Egyptian Arabic absorbed linguistic elements from Greek, Turkish, Italian, French, and English during various periods of occupation and cultural exchange. Regional variations within Egypt further enrich its linguistic diversity: northern regions, including the Delta and Cairo, feature subtle dialectal differences, while Upper Egypt’s Sa’idi Arabic retains more conservative features. Additionally, Bedouin communities in the Western Desert speak Arabic varieties that are distinct from urban Egyptian Arabic Souag (2009); Magidow (2021).
@@ 2.2 Participants
All participants, including both recruits and their peers, signed consent forms explicitly detailing the use and processing of their data in accordance with Swiss law. After the corpus was collected, it was manually anonymized to ensure privacy, and all real names were removed and substituted with fictitious ones where necessary. Subsequently, a professional Arabic-speaking translator translated the corpus into MSA and English, with these translations serving as reference texts for automatic metrics. The translation into Arabic represents an intralingual transformation from a dialectal and informal variety of Arabic to a formal and standardized form.
Table 1 presents the collected corpora, including the number of segments and tokens, the average number of tokens per sentence and the percentage of foreign and mixed words (code-mixing). Mixed words are created by combining roots from one language with prefixes, suffixes, or morphological patterns from another language, reflecting linguistic creativity and contact-induced change.
@@ 3 Experimental Setup
  • model $\in$ [GPT-4o, Llama 3, Claude, Gemini, Gemma, Mistral, Jais].
All models were prompted with a temperature of $0.5$. A discussion of the chosen models and prompts is available in 3.1 and in 3.2. Evaluation metrics are presented in 3.3. Our code is available. [^3] https://github.com/iguanodon-ai/ArabizivsLLMs
@@ 3.1 Models
The models used in the experiments are all decoder-only transformer Radford et al. (2018) models generally called “generative LLMs”. We used a range of instruction-tuned LLMs of different parameter sizes (from 27B for Gemma to at least 70B for Llama3, while proprietary models are expected to be much larger) to cover various models, from open weights to proprietary, general purpose or, in the case of Jais, ones that specifically target the English-Arabic pair Sengupta et al. (2023).
In this paper, “Llama” refers Llama 3.3 70B-Instruct (Dubey et al., 2024), “GPT-4o” (OpenAI, 2024) is gpt-4o-2024-08-06, [^4] https://arxiv.org/pdf/2203.02155 “Claude” to Anthropic ’s 3.5 Sonnet,555anthropic.claude-3-5-sonnet-20240620-v1:0 “Gemma” to gemma-2-27b-it Gemma Team et al. (2024), and, also from Google, “Gemini” to the latest Gemini 1.5 Pro version Gemini Team et al. (2024). [^6] December 2024 release. “Mistral” is Mistral Large 24.11 from the eponymous company and, finally, “Jais” refers to jais-family-30b-16k-chat Sengupta et al. (2023).
@@ 3.2 Prompts
In order to achieve the best translation results, we built on He (2024)’s findings by assigning the role of a professional translator to our LLM. This approach outperformed both simpler prompts and those with excessive context. Furthermore, for each of the three main dialects, we used three prompt strategies: no-shot, one-shot, and two-shot, all written in English. These prompts were the same across regions, except for the specific mention of each dialect in the corresponding prompts. The examples used in the one- and two-shot configurations are not part of the evaluated set, and are from the Algerian and Lebanese dialects. We further refined and duplicated these prompt variations to cover two target languages: one set asked for translation into English and the other into MSA. Finally, all prompts were translated into Arabic by a native Arabic speaker who is also a professional translator. In total, we ran experiments with 36 unique prompts (3 regions * 3 strategies * 2 target languages * 2 prompt languages), or 18 per target language, which we used on all models. The prompts in English are available in Appendix A, while their equivalents in Arabic can be found here.
@@ 3.3.2 Human Evaluation
For human evaluation, we adopted the direct assessment method, which evaluates translations based on a Likert scale ranging from 1 to 5 (higher is better) according to two key criteria: fluency and adequacy (See Appendix B). Due to time and human resource constraints, we did not manually annotate all translations. We instead sampled a random machine translation for both target languages and for each source sentence of our dataset. These machine translation outputs were sampled across all our variables, i.e. models, prompt languages, and prompt strategies. The resulting set, consisting of 671 segments (268 for Lebanon, 159 for Egypt and 244 for Algeria, see Table 1) for each target language, was then manually rated by two native speakers of Arabic who are professional translators, one of them being the first author of this study and the original translator of the dataset. We then calculated Cohen (1960) $\kappa$ to measure their agreement in terms of fluency and accuracy (see Appendix E for results). Cohen (1960) $\kappa$ results indicate moderate agreement for adequacy and lower agreement for fluency, with some variations across language pairs.
The set, which not only consisted of the reference and machine translation but also of the source sentence, was then iteratively fed into GPT-4o in an “LLM-as-a-judge” setting (Zheng et al., 2023), with a prompt in English tasking the LLM to follow the human annotation guidelines. [^8] The prompt is shared in Appendix C. The LLM showed strong correlation with both human annotators, with Spearman (1904)’s $\rho$s comprised in the range from $0.457$ (annotator 2, fluency, Egypt to EN) to $0.844$ (annotator 2, adequacy, Egypt to AR). These results indicate that an LLM can be used as an easy way to gauge translation quality during model development. The different correlations as well as all the data for direct assessment can be found in Appendix E.
@@ 4.1 Qualitative Error Analysis
Due to space constraints, this section will only provide some examples of the main errors. Refer to the table for a more detailed overview of the main errors. Most models tend to mistranslate, especially when figurative language is used. A larger issue lies with Llama3 which tends to output words in another script when translating to MSA. An obvious example is the translation of the segment Almatar da (\<هذا المطار>, “this airport”) as \<المطر> да – transforming the “da” in Arabic to Cyrillic. The problem is not limited to Cyrillic, as characters in Latin and Chinese scripts can also be found in the output. Another type of failure specific to a model is Jais’ re-occuring hallucinations. The model often associates feelings of anger to an otherwise neutral message, leading to translations that are irrelevant and contain violent information.
@@ 6 Limitations
7 Acknowledgements
We would like to thank the volunteers who donated their conversations, the translator, the reviewers for their helpful comments and suggestions. Finally, our thanks go to Dr. Thien for her proofreading assistance.
@@ Appendix D Metrics Signatures
BLEU: nrefs:1|case:mixed|eff:no|tok:13a|smooth:exp|version:2.5.1
chrF: nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.5.1
TER: nrefs:1|case:lc|tok:tercom|norm:no|punct:yes|
        asian:no|version:2.5.1
@@ Appendix F Automatic Evaluation Results for all Models
Due to space constraints, we are presenting the results averaged over the two prompt languages. A more comprehensive overview of the results is available in the accompanying spreadsheet.
@@
GPT-4o EN no-shot 17.386 43.081 77.038 0.478 0.434 0.733
GPT-4o EN one-shot 20.158 45.232 70.287 0.529 0.439 0.757
GPT-4o EN two-shot 19.771 45.496 70.294 0.521 0.439 0.755
GPT-4o AR no-shot 8.395 36.637 86.501 0.558 0.422 0.757
GPT-4o AR one-shot 10.099 38.221 78.701 0.586 0.423 0.776
GPT-4o AR two-shot 10.150 38.090 79.775 0.585 0.421 0.774